### v1

# VCNet: Recreating High-Level Visual Cortex Principles for Robust Artificial Vision

Abstract
Despite their success in image classification, modern convolutional neural networks (CNNs) exhibit fundamental limitations, including data inefficiency, poor out-of-distribution generalization, and vulnerability to adversarial perturbations. The primate visual system, in contrast, demonstrates superior efficiency and robustness, suggesting that its architectural principles may offer a blueprint for more capable artificial vision systems. This paper introduces Visual Cortex Network (VCNet), a novel neural network architecture whose design is informed by the macro-scale organization of the primate visual cortex. VCNet emulates key biological mechanisms, including hierarchical processing across distinct cortical areas, dual-stream information segregation, and top-down predictive feedback. We evaluate VCNet on two specialized benchmarks: the Spots-10 animal pattern dataset and a light field image classification task. Our results show that VCNet achieves a classification accuracy of 92.1% on Spots-10 and 74.4% on the light field dataset, surpassing contemporary models of comparable size. This work demonstrates that integrating neuroscientific principles into network design can lead to more efficient and robust models, providing a promising direction for addressing long-standing challenges in machine learning.
1 Introduction
Contemporary deep learning models for image recognition, while powerful, face critical challenges that impede their widespread deployment. These models often require extensive labeled training data Krizhevsky et al. (2012), exhibit poor generalization to out-of-distribution examples Sagawa et al. (2020), and are notoriously vulnerable to adversarial attacks and partial occlusion Liu et al. (2022). Minor, human-imperceptible perturbations or hidden object parts can cause catastrophic failures in prediction, raising concerns about their reliability in safety-critical applications. Furthermore, the escalating computational and energy costs associated with training state-of-the-art models present significant barriers to research and development Tan and Le (2019). These persistent issues motivate a re-evaluation of the prevailing architectural paradigms in computer vision.
In contrast, the primate visual system is a paragon of efficiency and robustness. Humans can learn to recognize objects from few examples Lake et al. (2015), generalize effortlessly across novel contexts Geirhos et al. (2018), robustly identify occluded objects Hegdé et al. (2008), and operate with unparalleled energy efficiency Lennie (2003). These capabilities are rooted in the specific architectural and computational principles of the visual cortex, notably its hierarchical organization Felleman and van Essen (1991); Grill-Spector and Malach (2004) and its use of predictive processing Rao and Ballard (1999); de Lange et al. (2018).
In this work, we explore this biological paradigm by proposing VCNet, a novel neural network whose macro-architecture is derived from the primate visual cortex. We systematically incorporate principles of the brain’s visual pathways to develop a model that aims to be more robust and efficient. Our contributions are threefold:
  • We introduce VCNet, a deep neural network architecture that models the high-level information flow between major areas of the visual cortex, including dual-stream processing, recurrent connections, and top-down predictive feedback.
  • We demonstrate the efficacy of VCNet on the Spots-10 animal pattern benchmark, selected to test our bio-inspired architecture on a task that mirrors a key evolutionary pressure for vision, and show that it outperforms other models of comparable size.
  • We further evaluate VCNet on a light field image classification task, providing evidence that its bio-inspired design is particularly well-suited for processing richer, multi-view data that more closely approximates the input to the human visual system.
2 The VCNet Architecture
While a complete replication of the visual system is infeasible, our research focuses on emulating the macro-scale organization of the visual cortex, including the connectivity patterns and relative computational capacity of its distinct regions. This structure serves as a scaffold for integrating more detailed biological computations. VCNet is a deep neural architecture engineered to operationalize these principles.
2.1 Biologically-Inspired Design Principles
Our model’s design is predicated on two foundational principles of primate vision: its hierarchical organization and its reliance on predictive feedback.
Hierarchical Processing in the Visual Cortex
Visual information propagates from the retina through a hierarchy of cortical areas (V1, V2, V3, V4, V5), each specialized for extracting progressively complex features Huff et al. (2023a). The primary visual cortex (V1) detects simple elements like oriented edges. It projects to V2, which processes intermediate features like contours and color. V2, in turn, projects to higher-order areas: V4, which is crucial for color and form perception, and V5 (or MT), which is specialized for motion Huff et al. (2023b). This intricate connectivity, mapped using neuronal tracing techniques Fulton (2001), forms a highly efficient cascade of feature extractors Sheth and Young (2016), as illustrated in Figure 1.
Figure 1: A high-level model of information pathways in the primate visual cortex, illustrating the hierarchical series of feature extraction stages Fulton (2001). This organization forms the architectural basis of VCNet.
Predictive Coding
The visual cortex is not a purely feedforward system. It employs predictive coding, a mechanism where higher-level cortical areas send top-down predictions of expected sensory input to lower-level areas. The bottom-up signals carry the actual sensory information, and discrepancies between predictions and inputs generate prediction errors. These error signals are propagated up the hierarchy to update and refine the brain’s internal model of the world, thereby minimizing future prediction errors Lowet and Uchida (2024); Urgen and Miller (2015).
2.2 Architectural Framework
Departing from conventional, monolithic CNN architectures, VCNet is structured as a directed acyclic graph that models the known connectivity between the major visual cortical areas. The channel capacity of each module is scaled to approximate the relative neuronal populations in its biological counterpart. The architecture is built around the two primary visual processing pathways.
Ventral Stream
This "what" pathway models object recognition, progressing from V1 through modules representing V2 (Interstripe, Thin Stripe), V4, and the inferotemporal (PIT, CIT, AIT) cortices. It is specialized for extracting features related to form and identity.
Dorsal Stream
This "where/how" pathway models spatial and motion analysis, flowing from V1 through V2 (Thick Stripe), the middle temporal (MT) and medial superior temporal (MST) areas, and onward to parietal regions.
These streams are interconnected at multiple levels, enabling the integration of object identity with spatial information. The final representation is formed in the AIT module, which receives convergent inputs and feeds into the classification layer. VCNet’s functionality is realized through several specialized computational blocks.
2.2.1 Multi-Scale Feature Extraction (V1)
To emulate the diverse receptive field sizes in the primary visual cortex, the V1 module processes input through three parallel depthwise separable convolution streams with different kernel sizes (3x3, 5x5, 7x7). The resulting feature maps are concatenated, providing a rich, multi-scale representation to all subsequent layers.
2.2.2 Recurrent Processing Blocks (MT/MST)
To model the iterative refinement of representations observed in cortical computation, the MT and MST modules for motion processing incorporate Recurrent Blocks. These blocks apply a convolutional transformation with shared weights for a fixed number of iterations ($t=3$), with each iteration receiving the output of the previous one plus a residual connection from the initial input.
2.2.3 Attentional Modulation (CBAM)
To emulate the brain’s ability to focus on salient features, key modules (V1, MT, V4) incorporate a Convolutional Block Attention Module (CBAM). CBAM sequentially infers and applies channel-wise and spatial attention maps, allowing the network to adaptively reweight and select the most informative features.
2.2.4 Lateral Interaction Module (V1)
The V1 module includes a Lateral Interaction block, implemented as a convolution followed by channel-wise self-attention within a residual connection. This simulates the horizontal connections within cortical layers that mediate contextual effects like lateral inhibition, crucial for edge enhancement.
2.2.5 Predictive Coding Loop
We implement predictive coding via a top-down connection from the highest level of the ventral stream (AIT) back to V1. The AIT module generates a prediction of V1 feature activations. This prediction is subtracted from the actual bottom-up V1 activity to compute a prediction error, $\epsilon=\text{ReLU}(\text{V1}_{\text{bottom-up}}-\text{AIT}_{\text{top-down}})$. This error signal serves as a potent learning signal, driving the network to refine its internal representations.
2.2.6 Neuromodulatory Gating
To model the global gain control exerted by neuromodulators, we introduce a ‘Neuromodulation‘ block in key modules (V1, MT, V4). This block applies a learnable, channel-wise multiplicative scaling factor to feature maps, allowing the network to dynamically adjust the excitability of different feature pathways.
3 Experiments and Results
We benchmarked VCNet’s performance against contemporary neural networks of comparable size to assess its image classification capabilities, focusing on data modalities that are particularly relevant to the evolution and function of biological vision.
3.1 Experiment 1: Animal Pattern Classification
Motivation
Key evolutionary drivers for primate vision include finding food and avoiding predators, tasks that rely heavily on pattern recognition Kaas (2012); Fornalé et al. (2012). The primate visual cortex is thus highly optimized for this purpose. We therefore evaluated our biologically-inspired model on a benchmark focused on classifying animal patterns.
Methodology
We utilized the Spots-10 dataset, which contains 50,000 grayscale 32x32 pixel images across 10 classes of animal patterns Atanbori (2024). We trained VCNet and compared its performance against a suite of established models.
Table 1: Test accuracy and model size on the Spots-10 benchmark. Best values are in bold.
Model Test Accuracy (%) Model Size (MB)
VCNet Mini (Ours) 92.08 0.04
DenseNet121 Distiller 81.84 0.07
ResNet101V2 Distiller 80.29 0.07
ResNet50V2 Distiller 79.03 0.07
MobileNet Distiller 78.26 0.07
MobileNetV3-Small Distiller 78.04 0.07
Results
As shown in Table 1, VCNet Mini attains the highest accuracy on Spots-10 (92.08%), outperforming the strongest baseline (DenseNet121 Distiller, 81.84%) by 10.24 percentage points. To ensure a fair comparison with the lightweight distilled baselines, we reduced VCNet’s hidden-layer widths to form the Mini variant. VCNet Mini uses only 0.04 MB of storage, about 43% smaller than the 0.07 MB baselines—while delivering the best accuracy. These findings indicate that architectures inspired by visual-cortex information flow can yield models that are both highly accurate and extremely compact on this benchmark.
3.2 Experiment 2: Light Field Classification
Motivation
Standard 2D images are flat projections of the 3D world, discarding vast amounts of visual information. The human visual system (HVS) processes a much richer input, leveraging binocular vision and eye movements to interpret a subset of the 7D plenoptic function Adelson and Bergen (1991). This allows it to perceive a robust 3D representation of a scene by using cues from the light field, such as parallax and view-dependent reflectance Xia et al. (2014). Light field cameras, which capture both the intensity and the angular direction of light rays, provide data that is a much closer analogue to the input processed by the HVS Lin et al. (2024). We hypothesize that an architecture designed to emulate the visual cortex will demonstrate superior performance when provided with input data that more closely matches the richness of biological vision.
Methodology
We evaluated VCNet on a light field image classification task using a standard dataset Raj et al. (2016) and compared its performance against benchmark models: ResNet18, VGG11 with Batch Normalization, and MobileNetV2.
Table 2: Performance and Size Comparison on Light Field Image Classification.
Model Test Accuracy (%) Model Size (MB)
VCNet (Ours) 74.42 3.52
MobileNetV2 72.09 8.66
ResNet18 65.12 42.69
VGG11_BN 51.16 491.39
Results
The results, summarized in Table 2, highlight VCNet’s superior performance and efficiency. VCNet achieved the highest test accuracy (74.42%) while maintaining a minimal model size of 3.52 MB. This is over ten times smaller than ResNet18 and over 100 times smaller than VGG11. This result highlights the efficacy of VCNet’s bio-inspired design for processing high-dimensional visual data, validating our architectural choices.
4 Conclusion and Future Work
In this work, we introduced VCNet, a novel architecture whose design is guided by the computational principles and anatomical organization of the primate visual cortex. By incorporating mechanisms such as hierarchical dual-stream processing, recurrence, and predictive coding, VCNet demonstrates superior performance and parameter efficiency on specialized image classification tasks compared to conventional CNNs. Our findings underscore the significant potential of neuroscience-inspired AI to address fundamental challenges in machine learning. This convergence of disciplines not only offers a path toward more capable artificial systems but also provides computational frameworks for testing hypotheses about brain function.
Our model opens several avenues for future research. One direction is to develop more specialized modules for each cortical area (V1, MT, V4), allowing for finer-grained architectural exploration and ablation studies. Further investigation into more biologically plausible mechanisms, such as alternative activation functions or more sophisticated predictive coding schemes with precision-weighting and temporal prediction, could enhance model robustness. Finally, integrating reinforcement learning could allow the model to learn adaptive visual representations tied to behavioral goals, potentially offering a principled solution to the challenge of out-of-distribution generalization.
Author Contributions
Brennen Hill: Project lead, conceptualization, software, engineering, investigation, research, writing
Zhang Xinyu: Software, engineering, investigation, research, writing.
Timothy Putra Prasetio: Software, engineering, investigation, research, writing.